An electrical firm manufactures light bulbs that have a length of life that is normally distributed with mean equal to 799 hours and a standard deviation of 49 hours. Find the probability that a bulb burns between 769 and 878 hours.


1
Expert's answer
2022-02-01T15:08:27-0500

We use Z calculations

Z(769)= (769 - 799) / 49 = -0.61224898

Z(878)=(878 - 799) / 49 = 1.612244898

P(769 < X < 878)=P (-0.61 < Z < 1.61)

From the normal table we obtain

(0.94630 - 0.27093) = 0.67537

Thus our answer is 0.67537